OpenStudy (anonymous):

Algebra 2 help please??

OpenStudy (phi):

can you re-write your equation so it is in standard form ?

OpenStudy (anonymous):

probably..? how do you do that though?

OpenStudy (phi):

standard form is the form you want so it matches the quadratic formula ax^2 + bx + c =0

OpenStudy (anonymous):

so what would my equation look like?

OpenStudy (anonymous):

It would look like it is in standard form once you place it in standard the form of: \[ax^2 + bx +c =0 \] instead of \[ax^2 +bx = c\]

OpenStudy (anonymous):

You can't start solving it until you do that 1st step.